# Toward a Brazilian Research Agenda in Quantum Software Engineering: A Systematic Mapping Study

### v1

@@ Abstract.
Context: Quantum Software Engineering (QSE) has emerged as a key field to support the development of reliable, maintainable, and scalable quantum applications, bridging advances in quantum computing with established practices in software engineering. Problem: Despite its growth, the field still suffers from fragmented knowledge, with a lack of standardized methodologies, tools, and guidelines tailored to the unique features of the quantum paradigm. Additionally, countries like Brazil have had limited participation in the development of this emerging domain. Objective: This study aims to map the state of the art in QSE by identifying current research trends, recurring contributions, and existing gaps that can guide future investigations and strategic initiatives. Methodology: A systematic mapping study was conducted analyzing selected publications based on inclusion and exclusion criteria. Articles were categorized by study type, research type, and alignment with the SWEBOK knowledge areas. Results: Most of the reviewed studies are primary research articles written in English, with a strong focus on Software Engineering Models and Methods, Software Architecture, and Software Testing. Conceptual proposals and technical solutions predominate, while empirical validations remain limited. Conclusions: Findings confirm that QSE is a promising but still maturing field. The standardization of practices, expansion of empirical studies, and inclusion of researchers from developing countries are crucial for advancing the discipline. Additionally, Brazilian contributions are still scarce, highlighting the urgent need to establish a national research agenda. As a main contribution, this study proposes a Brazilian Research Agenda in QSE, outlining priority areas and opportunities to foster a local scientific community and accelerate progress in this emerging field.
@@
conference: Brazilian Symposium on Software Engineering; September 22–26, 2025; Recife, PEcopyright: none
@@ 1. Introduction
In this context, the emerging field of Quantum Software Engineering (QSE) has gained attention as a research frontier that aims to systematize the development, testing, debugging, and maintenance of quantum software (Ali et al., 2022). Additionally, QSE landscape remains fragmented and lacks a consolidated research agenda, especially in regions like Latin America (Álvaro M. Aparicio-Morales et al., 2024). Therefore, there is a pressing need to understand the current state of QSE research.
@@ 3.1. Planning
  • RQ5: Which types of studies and research strategies are most prevalent?
  • RQ6: What areas of SWEBOK are most frequently addressed in QSE research?
@@
In this study, ACM Digital Library, IEEE Digital Library, Scopus, and Web of Science databases were selected due to extensive coverage of peer-reviewed literature across various disciplines, including computer science, engineering, and social sciences.
To ensure a comprehensive search, the search string was formulated based on the PICO framework, focusing on the Population and Intervention elements for being a SMS, as recommended by (Kitchenham and Charters, 2007). Considering QSE an adaptation of traditional Software Engineering (SE) based on QC fundamentals, we defined QC synonymous representing the population and Software Engineering Body of Knowledge (SWEBOK) areas as intervention related to SE. This approach was employed to retrieve studies addressing QSE context. The final search string is presented in Table 1.
@@
The search was performed by applying the search string to the metadata (i.e., title, abstract, and keywords) of each article across the source. Inclusion and exclusion criteria were defined to filter publications relevant to the study. Therefore, the inclusion criteria are (i) article must be in the QSE context, (ii) article must be from the latest research, and (iii) article published in conference, journal or book chapters. Exclusion criteria are (i) articles that cannot be accessed completely and (ii) articles that do not address aspects related to software production for the quantum computing paradigm (e.g., quantum computing algorithms, quantum mechanics, quantum information).
@@ 3.2. Conducting
The application of the review protocol yielded the following results (see Table 2). Between February and April 2025, 3294 results were found in the databases. From these articles, the screening process began. Following inclusion and exclusion criteria, 3106 studies were excluded, resulting in 188 articles for data extraction. Parsifal [^3] https://parsif.al was used to support the screening process, Microsoft Excel to tabulate and export the data to be used in the analysis and generation of graphs in Google Colab [^4] https://colab.google/.
Due to the number of selected articles, we have made supplementary material available [^5] https://github.com/filipefernandesphd/SBES2025 (e.g., full list of selected studies, scripts used to generate graphs, among others) that can be used to aggregate in the analysis of the main results described in Section 4 and Section 5.
Source #papers #papers #final
selected excluded papers
@@
Scopus 1435 127 158
Web of Science 136 1707 29
Total 3294 3106 188
@@ 4.1. What is the temporal distribution of QSE publications? (RQ1)
To answer the first research question of this study, we analyzed the evolution in the number of studies related to QSE published over the years. According to Figure 3, we identified three key periods of publications. In the first period, the amount of publications are not relevant between 2002 and 2017, indicating that QSE research was incipient. In the second period, the publications started a moderate growing between 2018 and 2020. Finally, in the third period, we can see an expressive increase in publications from one year to the next, i.e., each year from 2021, the number of works in QSE has increased, arriving in 2024 with the historical maximum of 55 publications.
As presented in Section 3, the search for paper was performed between February and April 2025, which justifies the number of papers published. However, the number of publications in this period exceeds 2020, indicating an increasing number of works by the end of 2025.
@@ 4.2. Who are the most active institutions and authors in the field? (RQ2)
It is noted that the publication indicators of the institutions are not directly influenced by their associated researchers. This is the case of researcher Shaukat Ali (17), who stands out as the second most influential researcher in QSE. However, the institution to which he is associated (Simula Reseach Laboratory) is fourth in the ranking of institutions, with 29 publications.
@@ 4.3. Which countries are leading QSE research? (RQ3)
Table 3 complements Figure 6 by showing the 20 countries with the highest number of publications, grouped by continent. The leader is Spain (43), followed by Japan and Finland, both with 23 publications. Next come Germany (22), China (19), Italy (18), United States (18), Norway (14) and Canada (14). This indicates that most of the production is concentrated in developed countries.
When grouping countries by continent, it can be seen that Europe is the main hub of contributions in QSE, bringing together more than half of the countries listed. Asia also demonstrates relevance. In the Americas, although the United States is among the most productive, countries such as Brazil, Argentina and Chile appear with low participation. Oceania is represented only by Australia, with 11 publications.
@@
Europe Spain 43
Finland 23
Germany 22
Italy 18
Norway 14
@@
China 19
India 7
Saudi Arabia 6
@@
Argentina 4
Brazil 3
Chile 2
Oceania Australia 11
@@ 4.4. What are the most frequent publication venues? (RQ4)
From our analysis, we identified that most of the studies were published in conferences, with emphasis on the International Conference on Quantum Computing and Engineering (16), followed by the International Conference on Quantum Software, International Workshop on Quantum Software Engineering, International Workshop on Quantum Software Engineering and Technology, all conferences with 8 published papers. Information and Software Technology is the scientific journal that stands out from the others with 9 publications.
In general, venues such as Transactions on Software Engineering and Information and Software Technology are well-established journals, indicating that part of the research in QSE is being absorbed by traditional SE literature. In contrast, specialized workshops in the area have been created from conferences in SE and related fields to address QSE and encourage collaboration among researchers, as it is the case of the International Workshop on Quantum Software Engineering, International Workshop on Quantum Software Engineering and Technology, and International Workshop on Quantum Programming for Software Engineering.
@@
4.5. Which types of studies and research strategies are most prevalent? (RQ5)
The aim of this research question is to analyze the types of studies found, as well as the research strategies adopted in the primary studies. We classified the studies according to the types defined by (Kitchenham and Charters, 2007): primary study as an empirical study investigating a specific research question; secondary study as a study that reviews all the primary studies relating to a specific research question with the aim of synthesizing evidence related to a specific research question; and finally tertiary study as a review of secondary studies related to the same research question.
Table 4 shows that 88.8% of the identified papers are primary studies, while only 11.2% correspond to secondary studies. We expected that most of the articles would be primary studies. Unlike most systematic mappings, we decided to include secondary and tertiary studies in our work because we wanted to observe the number of contributions of this type of study in relation to the primary ones. We understand that 21 secondary studies is a very expressive number of publications, considering that QSE is a very recent research area.
Considering this scenario, it becomes pertinent to investigate the predominant methodological approaches among primary studies. These studies were classified according to the six well-established research types proposed by (Wieringa et al., 2006), namely: evaluation research, proposal of solution, validation research, philosophical papers, opinion papers, and personal experience papers.
@@ 4.5. Which types of studies and research strategies are most prevalent? (RQ5)
Table 4. Distribution of Study Types
Study Type Frequency %
Primary Study 167 88,8%
Secondary Study 21 11,2%
Total 188 100%
@@
4.6. What areas of SWEBOK are most frequently addressed in QSE research? (RQ6)
@@ 5.2. SWEBOK Focus by Research Strategy
Furthermore, Evaluation research is also significant in Software Testing, Software Architecture and Software Quality. In this type of research, there is a clearer concern with the empirical evaluation of techniques, tools or models applied to real or simulated contexts. The significant presence of this type of study in these categories may indicate a greater degree of methodological maturity, with an emphasis on verifying the practical effectiveness of the proposed solutions, which contributes to strengthening the body of knowledge of QSE based on concrete evidence. The prominence of these areas can be understood in light of the fundamental challenges imposed by the quantum paradigm on the software development process, such as probabilities, non-determinism, entanglement and superposition of states.
@@ 5.3. Geographic and Institutional
According to Figure 13, most institutions with significant production are located in Europe (Spain, Finland, Norway, Germany, Austria, Italy, Portugal). This may indicate that European research centers are mobilizing coordinated efforts to position themselves as leaders in QSE. There is strong funding from European programs such as Horizon Europe [^6] https://research-and-innovation.ec.europa.eu, in addition to international collaborations within the European Union that facilitate scientific exchange and shared funding.
@@
Spain stands out in particular with three highly productive institutions, namely University of Castilla-La Mancha (24) and University of Extremadura (14) universities; and aQuantum (7) as a company specialized in research, development, consulting and services in the fields of QSE and development. Spain is not traditionally a global leader in quantum computing, but the data suggests the existence of groups strongly dedicated to QSE, such as Mario Piattini, Ricardo Pérez-Castillo and Enrique Moguel, as shown in (Figure 5).
@@ 6. Brazilian Research Agenda
Given the gaps identified in this study, we propose strategies that can direct research and guide the strengthening of the Brazilian community in QSE.
@@ 7. Conclusion
Data used in the analysis, as well as supplementary materials can be accessed in https://github.com/filipefernandesphd/SBES2025.

### v2

@@ Abstract.
Context: Quantum Software Engineering (QSE) has emerged as a promising discipline to support the development of quantum applications by integrating quantum computing principles with established software engineering practices. Problem: Despite recent growth, QSE still lacks standardized methodologies, tools, and guidelines, and countries like Brazil have minimal representation in its development. Objective: This study aims to map the current state of QSE, identifying research trends, contributions, and gaps to support future investigations and initiatives. Methodology: A systematic mapping study was conducted across major databases, retrieving 3219  studies, of which 3052  were excluded, resulting in 167  studies for analysis. Publications were classified by study type, research type, and SWEBOK alignment. Results: The majority studies focused on Software Engineering Models and Methods, Software Architecture, and Software Testing. Conceptual and technical proposals dominate, while empirical validations are still few. Conclusions: QSE remains a maturing field. Standardization, empirical research, and the inclusion of developing countries are essential. As a key contribution, this study proposes a Brazilian Research Agenda in QSE to guide national efforts and foster a strong scientific community.
@@
conference: International Workshop on Quantum Software Engineering and Technology; Aug 31 –- Sep 5, 2025; USAcopyright: none
@@ 1. Introduction
In this context, the emerging field of Quantum Software Engineering (QSE) has gained attention as a research frontier that aims to systematize the development, testing, debugging, and maintenance of quantum software (Ali et al., 2022). Additionally, QSE landscape remains fragmented and lacks a consolidated research agenda, especially in regions like Latin America (Álvaro M. Aparicio-Morales et al., 2024a). Therefore, there is a pressing need to understand the current state of QSE research.
@@ 2.2. Development of Quantum Software
3. Related Work
Several secondary studies have explored the field of Quantum Software Engineering (QSE) from different perspectives. From De Stefano et al. (2024) have a systematic mapping on recurring topics and technologies used, while Khan et al. (2023) perform a systematic review aimed at software architecture for quantum systems. Serrano et al. (2022) and Dwivedi et al. (2024) offer panoramic views on components, tools and the life cycle of quantum software development, and Hacaloglu et al. (2024) discuss the functional requirements and their measurement in the quantum context. Other studies, such as Álvaro M. Aparicio-Morales et al. (2024b), focus on the characterization of the Latin American scenario, with emphasis on collaborations, regional challenges and priority research lines. (Piattini et al., 2024) and Dwivedi et al. (2024) also discuss trends and gaps in QSE practice, based on broad surveys about the state of the art.
Despite the relevant contributions of these works, none of them focus on the proposition of a national research agenda, articulating specific strategic axes for the coordinated advance of QSE in a country with low representativeness, as is the case of Brazil. This is the main differential of this study, which not only maps the state of the art of QSE, but proposes a structured plan divided into four axes - education and training, interdisciplinary research groups, national scientific events and funding strategies (see Section 7) - to foster the construction of an engaged and productive scientific community in Brazilian soil.
@@ 4.1. Planning
  • RQ5: Which research strategies are most prevalent?
  • RQ6: Which areas of SWEBOK are most frequently addressed in QSE research?
@@
In this study, ACM Digital Library, IEEE Digital Library, Scopus, and Web of Science databases were selected due to extensive coverage of peer-reviewed literature. To ensure a comprehensive search, the search string was formulated based on the PICO framework, focusing on the Population and Intervention elements for being a SMS, as recommended by Kitchenham and Charters (2007). Considering QSE an adaptation of traditional Software Engineering (SE) based on QC fundamentals, we defined QC synonymous representing the population and Software Engineering Body of Knowledge (SWEBOK) areas as intervention related to SE. This approach was employed to retrieve studies addressing QSE context. The final search string is presented in Table 1.
@@
The search was performed by applying the search string to the metadata (i.e., title, abstract, and keywords) of each article across the source. Inclusion and exclusion criteria were defined to filter publications relevant to the study. Studies that met all three inclusion criteria were selected and those that met at least one exclusion criterion were rejected. Inclusion criteria are (i) study must be in the QSE context, (ii) publication must be from the latest study, and (iii) it published in conference, journal or book chapters. Exclusion criteria are (i) publications as secondary studies, (ii) publications that can not be accessed completely, and (iii) publications that do not address aspects related to software production for the quantum computing paradigm (e.g., quantum computing algorithms, quantum mechanics, quantum information).
@@ 4.2. Conducting
The application of the review protocol yielded the following results (see Table 2). Between February and April 2025, 3219  results were found in the databases. From these articles, the screening process began. Following inclusion and exclusion criteria while reading the title, abstract and full text, 3052  studies were excluded, resulting in 3052  articles for data extraction. Parsifal [^3] https://parsif.al was used to support the screening process, Microsoft Excel to tabulate and export the data to be used in the analysis and generation of graphs in Google Colab [^4] https://colab.google/.
Due to the number of selected articles, we have made supplementary material available [^5] https://bit.ly/BRAgendaQSE (e.g., full list of selected studies, scripts used to generate graphs, among others) that can be used to aggregate in the analysis of the main results described in Section 5 and Section 6.
Source #retrieved #excluded #final
studies studies studies
@@
Scopus 1715 1626 89
Web of Science 1381 1304 77
Total 3219 3052 167
@@ 5.1. What is the temporal distribution of QSE publications? (RQ1)
To answer the first research question of this study, we analyzed the evolution in the number of studies related to QSE published over the years. According to Figure 3, we identified three key periods of publications. In the first period, the amount of publications are not relevant between 2002 and 2017, indicating that QSE research was incipient. In the second period, the publications started a moderate growing between 2018 and 2020. Finally, in the third period, we can see an expressive increase in publications from one year to the next, i.e., each year from 2021, the number of works in QSE has increased, arriving in 2024 with the historical maximum of 43 publications.
As presented in Section 4, the search for studies was performed between February and April 2025, which justifies the number of papers published. However, the number of publications in this period exceeds 2020, indicating an increasing number of works by the end of 2025.
@@ 5.2. Who are the most active institutions and authors in the field? (RQ2)
It is noted that the publication indicators of the institutions are not directly influenced by their associated researchers. This is the case of researcher Shaukat Ali (17), who stands out as the second most influential researcher in QSE. However, the institution to which he is associated (Simula Reseach Laboratory) is third in the ranking of institutions, with 29 publications.
@@ 5.3. Which countries are leading QSE research? (RQ3)
Table 3 complements Figure 6 by showing the 20 countries with the highest number of publications, grouped by continent. The leader is Spain (36), followed by Japan (23), Finland (19), as well as China, Germany and United States, both with 18 publications. This indicates that most of the production is concentrated in developed countries.
When grouping countries by continent, it can be seen that Europe is the main hub of contributions in QSE, bringing together more than half of the countries listed. Asia also demonstrates relevance. In the Americas, although the United States and Canada are among the most productive, Brazil appears with low participation (2). Oceania is represented only by Australia, with 10 publications.
@@
Europe Spain 36
Finland 19
Germany 18
Italy 16
Norway 13
@@
China 18
India 5
Saudi Arabia 5
@@
Canada 12
Brazil 2
Oceania Australia 10
@@ 5.4. What are the most frequent publication venues? (RQ4)
From our analysis, we identified that most of the studies were published in conferences, with emphasis on the International Conference on Quantum Computing and Engineering (16), followed by the International Conference on Quantum Software, International Workshop on Quantum Software Engineering and Technology, International Workshop on Quantum Software Engineering, all conferences with 8 published papers. Information and Software Technology is the scientific journal that stands out from the others with 7 publications. In general, venues such as Information and Software Technology and Transactions on Software Engineering are well-established journals, indicating that part of the research in QSE is being absorbed by traditional SE literature.
@@
5.5. Which research strategies are most prevalent? (RQ5)
The aim of this research question is to investigate the predominant methodological approaches among primary studies. These studies were classified according to the six well-established research types proposed by Wieringa et al. (2006), namely: evaluation research, proposal of solution, validation research, philosophical papers, opinion papers, and personal experience papers.
@@ 5.5. Which research strategies are most prevalent? (RQ5)
5.6. Which areas of SWEBOK are most frequently addressed in QSE research? (RQ6)
@@ 6.2. SWEBOK Focus by Research Strategy
Furthermore, Evaluation research is also significant in Software Testing, Software Architecture, Software Construction, and Software Quality. In this type of research, there is a clearer concern with the empirical evaluation of techniques, tools or models applied to real or simulated contexts. The significant presence of this type of study in these categories may indicate a greater degree of methodological maturity, with an emphasis on verifying the practical effectiveness of the proposed solutions, which contributes to strengthening the body of knowledge of QSE based on concrete evidence. The prominence of these areas can be understood in light of the fundamental challenges imposed by the quantum paradigm on the software development process, such as probabilities, non-determinism, entanglement and superposition of states.
@@ 6.3. Geographic and Institutional
According to Figure 13, most institutions with significant production are located in Europe (Spain, Finland and Norway). This may indicate that European research centers are mobilizing coordinated efforts to position themselves as leaders in QSE. There is strong funding from European programs such as Horizon Europe [^6] https://research-and-innovation.ec.europa.eu, in addition to international collaborations within the European Union that facilitate scientific exchange and shared funding.
@@
Spain stands out in particular with three highly productive institutions, namely University of Castilla-La Mancha (19) and University of Extremadura (13), as well as aQuantum (6) as a company specialized in research, development, consulting and services in the fields of QSE and development. Spain is not traditionally a global leader in quantum computing, but the data suggests the existence of groups strongly dedicated to QSE, such as Mario Piattini, Ricardo Pérez-Castillo and Enrique Moguel, as shown in (Figure 5).
@@
6.4. Threats to Validity
This study is subject to some threats to validity. Definition of search string and exclusive categorization by SWEBOK area may have limited the scope of the captured studies. Although four consolidated bases have been used, other relevant sources Springer Link, Engineering Village or even arXiv have not been included, which may impact the representativeness of the results. Finally, the emerging nature of QSE implies that the data analyzed, collected until the first trimester of 2025, can rapidly become outdated in the face of the acceleration of publications in the area.
@@ 7. Brazilian Research Agenda
Despite the global growth of QSE, led by countries such as Spain, Japan, Finland, China, Germany and the United States, Brazil’s involvement in the field is still incipient, with only 2 identified publications. This scenario highlights the urgent need for structured actions that boost the national presence in the area. Therefore, this work proposes a Brazilian research agenda in QSE, with the objective of articulating training, collaboration and financing to stimulate scientific production and consolidate Brazil as a relevant actor in this strategic field. Agenda is organized on four main axes: education and training, fostering interdisciplinary research groups, establishing national events and funding strategies.
@@ 8. Conclusion
Data used in the analysis, as well as supplementary materials can be accessed in https://bit.ly/BRAgendaQSE.